# Simulating Polynomial-Time Nondeterministic Turing Machines via Nondeterministic Turing Machines

### v21

@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
By Lemma 2.7 (respectively, Lemma 2.8) stated above, we only need to discuss the single-tape nondeterministic Turing machines (respectively, single-tape ${\rm co}\mathcal{NP}$ machines). Thus, in the following context, by a nondeterministic Turing machine, we mean a single-tape nondeterministic Turing machine. Similarly, by a ${\rm co}\mathcal{NP}$ machine we mean a single-tape ${\rm co}\mathcal{NP}$ machine.
@@
so that the tuple $(M,k)$ should be the concatenation of the binary string representing $M$ itself followed by the order $10^{k}1$. Now the tuple $(M,k)$ is encoded as a binary string, which can be explained as an integer.
@@
where $k_{1},k_{2}\in\mathbb{N}$. Since the Cantor pairing function is invertible (see [A8]), it is a bijection between $\mathbb{N}\times\mathbb{N}$ and $\mathbb{N}$. Note that we can further define a bijection $\tau$ between $\mathbb{N}$ and $\mathbb{N}_{1}$ such as $\tau:k\mapsto k+1$ where $k\in\mathbb{N}$ and $k+1\in\mathbb{N}_{1}$, which shows that there exists a $(1,1)$ correspondence between $\mathbb{N}$ and $\mathbb{N}_{1}$. As we have shown that any polynomial-time nondeterministic Turing machine is an integer, we can place any polynomial-time nondeterministic Turing machine and its order of polynomial into a tuple $(M,k)$ and use the Cantor pairing function to map the tuple $(M,k)$ to an integer in $\mathbb{N}_{1}$. The reader can easily check that such a method gives an enumeration of the set of polynomial-time nondeterministic Turing machines.
@@
In the following context, we often use the notation $\langle M\rangle$ to denote the shortest binary string that represents the same $n^{k}+k$ time-bounded nondeterministic Turing machine $(M,k)$. We should further remark that the shortest binary string $\langle M\rangle$ that represents the same $n^{k}+k$ time-bounded nondeterministic Turing machine $(M,k)$ is by concatenating in any order the codes for each of the quintuples in its next-move function followed by the order $10^{k}1$ without additional $1$ prefixed to it, from which it is clear that $\langle M\rangle$ may not be unique (since the order of codes of the quintuples may be different), but the lengths of all different $\langle M\rangle$ are the same.
@@ 4. Simulation of All Polynomial-time Nondeterministic Turing Machines
In computational complexity, the most basic approach to show hierarchy theorems uses simulation and diagonalization, because the simulation and diagonalization technique [Tur37, HS65] is a standard method to prove lower bounds on uniform computing models (i.e., the Turing machine model); see for example [AHU74]. These techniques work well in deterministic time and space measures; see e.g., [For00, FS07]. However, they do not work for nondeterministic time, which is not known to be closed under complement at present, i.e., one of the main issues discussed in this paper; hence it is unclear how to define a nondeterministic machine that “does the opposite;” see e.g., [FS07].
For nondeterministic time, we can still do a simulation but can no longer negate the answer directly. In this case, we apply the lazy diagonalization to show the nondeterministic time hierarchy theorem; see e.g., [AB09, SFM78, Zak83]. It is worth noting that Fortnow [For11] developed a much more elegant and simple style of diagonalization to show nondeterministic time hierarchy; see also [FS17]. Generally, lazy diagonalization [Zak83, AB09, SFM78] is a clever application of the standard diagonalization technique [Tur37, HS65, AHU74]. The basic strategy of lazy diagonalization is that for any $n\leq j<m=n^{t(n)}$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $1^{j+1}$, accepting if $M_{i}$ accepts and rejecting if $M_{i}$ rejects. When $j=m$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $x$ deterministically, accepting if $M_{i}$ rejects and rejecting if $M_{i}$ accepts. Since $m=n^{t(n)}$, $M$ has enough time to perform the trivial exponential-time deterministic simulation of $M_{i}$ on input $1^{n}$; what we are doing is deferring the diagonalization step by linking $M$ and $M_{i}$ together on larger and larger inputs until $M$ has an input large enough that it can actually “do the opposite” deterministically; see e.g., [FS07] for details.
However, although the set of all ${\rm co}\mathcal{NP}$ machines is proved to be enumerable and the enumeration is $e$ given in Section 3, it is challenging (even impossible) to apply the standard diagonalization technique or the lazy-diagonalization technique from [Zak83, AB09, FS07, SFM78] to create a language that is not in ${\rm co}\mathcal{NP}$. Instead, our main goal of this section is to apply simulation techniques rather than diagonalization techniques and the particular relations between $L(M)$ accepted by a polynomial-time nondeterministic Turing machine $M$ and $\overline{L}(M)$ accepted by a ${\rm co}\mathcal{NP}$ machine $M$ to show that there is a language $L_{s}$ accepted by a universal nondeterministic Turing machine with the property that $L_{s}\notin{\rm co}\mathcal{NP}$.
@@
    $U$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}$ machine $M_{l}$ for some $l$ (by Definition 2.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic Turing machine, i.e., $x$ represents the trivial Turing machine with an empty next-move function), then GOTO $5$, else determine $t$, the number of tape symbols used by $M_{l}$; $s$, its number of states; and $k$, its order. The third tape of $U$ can be used as “scratch” memory to calculate $t$.
@@ 5. Proving That $L_{s}\in\mathcal{NP}$
steps (in other words, $U$ runs at most $n^{i+1}$ steps for any $i\in\mathbb{N}_{1}$ where $n$ is the length of input, see Theorem 5.1 above), we have
@@
Note further that for any fixed $i\in\mathbb{N}_{1}$, $L_{s}^{i}$ is the language accepted by the nondeterministic Turing machine $U$ within time $O(n^{i})$, i.e., at most $n^{i+1}$ steps, we thus obtain
@@ 6. Breaking the “Relativization Barrier”
A ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}$ with oracle $A$ is exactly a polynomial-time nondeterministic oracle Turing machine $M^{A}$ with oracle $A$ working under the condition of “there exists no accepting path,” and we define the language accepted by ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}$ to be $\overline{L}(M^{A})$. Namely, for any input $w$, $w\in\overline{L}(M^{A})$ if and only if polynomial-time nondeterministic oracle Turing machine $M^{A}$ rejects $w$.
@@
First, note that $\mathcal{P}^{O}=\mathcal{NP}^{O}$, thus we have ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Clearly, in this scenario, to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$ is equivalent to separating $\mathcal{P}^{O}$ from ${\rm co}\mathcal{P}^{O}$. Moreover, if the set of all ${\rm co}\mathcal{P}^{O}$ machines is not enumerable, then of course we can not apply simulation techniques. Thus, we assume that the set of ${\rm co}\mathcal{P}^{O}$ machines is enumerable, as the premise of the theorem states. Since we have assumed that the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, the rest of the proof is similar to Remark 4.2. Namely, the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}^{O}$ via a universal deterministic oracle Turing machine $U^{O}$, because a polynomial-time deterministic oracle Turing machine that works under the “there exists no accepting path” condition is still a polynomial-time deterministic oracle Turing machine. In short, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic oracle Turing machines are symmetrical. Specifically, for an arbitrary deterministic oracle Turing machine $M^{O}$, if $M^{O}$ accepts the input $w$, then there is only one accepting path, and if $M^{O}$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic oracle Turing machines. If we apply the same techniques, i.e., the simulation techniques, to create a language not in ${\rm co}\mathcal{P}^{O}$, then the universal deterministic oracle Turing machine $U^{O}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{O})\in\mathcal{P}^{O}$ (the proof is similar to that of Theorem 6.4 below). But, by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic oracle Turing machines, we clearly have that $L(U^{O})\in{\rm co}\mathcal{P}^{O}$. Thus, we are unable to claim that $L(U^{O})\notin{\rm co}\mathcal{P}^{O}$.
@@
    The simulation can be done within in time
@@
Since the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines is enumerable, and apart from it, we further have the above rational assumptions, i.e., the following: (1) The ${\rm co}\mathcal{NP}^{A}$ machine can be encoded to a string over $\{0,1\}$; (2) There exist universal nondeterministic oracle Turing machines with oracle $A$ that can simulate any other polynomial-time nondeterministic oracle Turing machine with oracle $A$, which is similar to the behavior of universal nondeterministic Turing machine $U$ constructed in Section 4; (3) The simulation can be done within time
@@
    $U^{A}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$ with oracle $A$ for some $j$ (by Definition 6.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic oracle Turing machine $M^{A}_{j}$ for some $j$, i.e., $x$ represents the trivial oracle Turing machine with an empty next-move function), then GOTO $6$, else determine $t$, the number of tape symbols used by $M^{A}_{j}$; $s$, its number of states; and $k$, its order.  [^7] We suppose that the order $10^{k}1$ of $M^{A}_{j}$ for some $k\in\mathbb{N}_{1}$ is also encoded into the binary string representing ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$, which is similar to Section 3. The third tape of $U^{A}$ can be used as “scratch” memory to calculate $t$.
@@
    By using its nondeterministic choices, $U^{A}$ moves as per the path given by $x$. The moves of $U^{A}$ are counted in binary in the block of tape $3$. If the counter on tape $3$ overflows, then $U^{A}$ halts. $U^{A}$ accepts $x$ if and only if there is at least one computation path from the start state of $U^{A}$ leading to the accept state and the total number of moves can not exceed
@@
We claim that $L^{A}_{s}\notin{\rm co}\mathcal{NP}^{A}$. Indeed, suppose for the sake of contradiction that $L^{A}_{s}$ is decided by some $n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{i}$. Then, we may assume that $M^{A}_{i}$ is a two-tape (with an input tape and an oracle tape) $T(n)=n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}^{A}$ machine. By Definition 6.4, we can regard it as an $n^{k}+k$ time-bounded nondeterministic oracle Turing machine with oracle $A$, and we can further let $U^{A}$ on input $1^{n-m}\langle M^{A}_{i}\rangle$ simulate $M^{A}_{i}$ on input $1^{n-m}\langle M^{A}_{i}\rangle$. Let $M^{A}_{i}$ have $s$ states and $t$ tape symbols, and the shortest length of $M^{A}_{i}$ is $m$, i.e.,
@@
If $\mathcal{P}^{A}\neq\mathcal{NP}^{A}$. Assume that the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines is enumerable. Further, under the above rational assumptions, we have
@@
Under some rational assumptions (i.e., the above rational Assumptions (1), (2), and (3)), and if $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, then the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines is not enumerable. Thereby, ordinary simulation techniques will generally not apply to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.
Now, we are naturally at a right point to present the proof of Theorem 1.3 as follows: Proof of Theorem 1.3. Note that the statement of Theorem 1.3 is precisely the same as Theorem 6.6; therefore, we have done the proof.

### v22

@@ 3. All Polynomial-time Nondeterministic Turing Machines Are Enumerable
By Lemma 2.7 (respectively, Lemma 2.8) stated above, we only need to discuss the single-tape nondeterministic Turing machines (respectively, single-tape ${\rm co}\mathcal{NP}$ machines). Thus, in the following context, by a polynomial-time nondeterministic Turing machine, we mean a single-tape polynomial-time nondeterministic Turing machine. Similarly, by a ${\rm co}\mathcal{NP}$ machine we mean a single-tape ${\rm co}\mathcal{NP}$ machine.
@@
so that the tuple $(M,k)$ should be the concatenation of the binary string representing $M$ itself followed by the order $10^{k}1$. Now the tuple $(M,k)$ is encoded as a binary string, which still can be explained as an integer.
@@
where $k_{1},k_{2}\in\mathbb{N}$. Since the Cantor pairing function is invertible (see [A8]), it is a bijection between $\mathbb{N}\times\mathbb{N}$ and $\mathbb{N}$. Note that we can further define a bijection $\tau$ between $\mathbb{N}$ and $\mathbb{N}_{1}$ such as $\tau:k\mapsto k+1$ where $k\in\mathbb{N}$ and $k+1\in\mathbb{N}_{1}$, which shows that there exists a $(1,1)$ correspondence between $\mathbb{N}$ and $\mathbb{N}_{1}$. As we have shown that any polynomial-time nondeterministic Turing machine is an integer, we can place any polynomial-time nondeterministic Turing machine and its order of polynomial into a tuple $(M,k)$ and use the Cantor pairing function to map the tuple $(M,k)$ to an integer in $\mathbb{N}_{1}$. As a result, the reader can easily check that such a method gives an enumeration of the set of polynomial-time nondeterministic Turing machines.
@@
In the following context, we often use the notation $\langle M\rangle$ to denote the shortest binary string that represents the same $n^{k}+k$ time-bounded nondeterministic Turing machine $(M,k)$. We should further remark that the shortest binary string $\langle M\rangle$ that represents the same $n^{k}+k$ time-bounded nondeterministic Turing machine $(M,k)$ is by concatenating in any order the codes for each of the quintuples in its next-move function followed by the order $10^{k}1$ without additional $1$ prefixed to it, from which it is clear that $\langle M\rangle$ may not be unique (since the order of codes of the quintuples may be different), but the lengths of all different $\langle M\rangle$ are the same. Simultaneously, the notation $\langle M\rangle$ also denotes the shortest binary string that represents the same $n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}$ machine.
@@ 4. Simulation of All Polynomial-time Nondeterministic Turing Machines
In the field of computational complexity, the most basic approach to show hierarchy theorems uses simulation and diagonalization, because the simulation and diagonalization technique [Tur37, HS65] is a standard method to prove lower bounds on uniform computing models (i.e., the Turing machine model); see for example [AHU74]. These techniques work well in deterministic time and space measures; see e.g., [For00, FS07]. However, they do not work for nondeterministic time, which is not known to be closed under complement at present, i.e., one of the main issues discussed in this paper; hence, it is unclear how to define a nondeterministic machine that “does the opposite;” see e.g., [FS07].
For nondeterministic time, we can still do a simulation but can no longer negate the answer directly. In this case, we apply the lazy diagonalization to show the nondeterministic time hierarchy theorem; see e.g., [AB09, SFM78, Žák83]. It is worth noting that Fortnow [For11] developed a much more elegant and simple style of diagonalization to show nondeterministic time hierarchy; see also [FS17]. Generally, lazy diagonalization [Žák83, AB09, SFM78] is a clever application of the standard diagonalization technique [Tur37, HS65, AHU74]. The basic strategy of lazy diagonalization is that for any $n\leq j<m=n^{t(n)}$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $1^{j+1}$, accepting if $M_{i}$ accepts and rejecting if $M_{i}$ rejects. When $j=m$, $M$ on input $1^{j}$ simulates $M_{i}$ on input $x$ deterministically, accepting if $M_{i}$ rejects and rejecting if $M_{i}$ accepts. Since $m=n^{t(n)}$, $M$ has enough time to perform the trivial exponential-time deterministic simulation of $M_{i}$ on input $1^{n}$; what we are doing is deferring the diagonalization step by linking $M$ and $M_{i}$ together on larger and larger inputs until $M$ has an input large enough that it can actually “do the opposite” deterministically; see e.g., [FS07] for details.
However, although the set of all ${\rm co}\mathcal{NP}$ machines is proved to be enumerable and the enumeration is $e$ given in Section 3, it is challenging (even impossible) to apply the standard diagonalization technique or the lazy-diagonalization technique from [Žák83, AB09, FS07, SFM78] to create a language that is not in ${\rm co}\mathcal{NP}$. Instead, our main idea for this section is to apply simulation techniques rather than diagonalization techniques and the particular relations between $L(M)$ accepted by a polynomial-time nondeterministic Turing machine $M$ and $\overline{L}(M)$ accepted by a ${\rm co}\mathcal{NP}$ machine $M$ to show that there is a language $L_{s}$ accepted by a universal nondeterministic Turing machine with the property that $L_{s}\notin{\rm co}\mathcal{NP}$.
@@
    $U$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}$ machine $M_{l}$ for some $l$ (by Definition 2.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic Turing machine, i.e., $x$ represents the trivial Turing machine with an empty next-move function), then GOTO $5$; else determine $t$, the number of tape symbols used by $M_{l}$; $s$, its number of states; and $k$, its order. The third tape of $U$ can be used as “scratch” memory to calculate $t$.
@@ 5. Proving That $L_{s}\in\mathcal{NP}$
steps (in other words, $U$ runs at most $n^{i+1}$ steps for any $i\in\mathbb{N}_{1}$ where $n$ is the length of input; see Theorem 5.1 above), we have
@@
Note further that for any fixed $i\in\mathbb{N}_{1}$, $L_{s}^{i}$ is the language accepted by the nondeterministic Turing machine $U$ within time $O(n^{i})$, i.e., at most $n^{i+1}$ steps; we thus obtain
@@ 6. Breaking the “Relativization Barrier”
A ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}$ with oracle $A$ is exactly a polynomial-time nondeterministic oracle Turing machine $M^{A}$ with oracle $A$ working under the condition of “there exists no accepting path,” and we define the language accepted by ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}$ to be $\overline{L}(M^{A})$. Namely, for any input $w$, $w\in\overline{L}(M^{A})$ if and only if polynomial-time nondeterministic oracle Turing machine $M^{A}$ rejects $w$ as the input to $M^{A}$.
@@
First, note that $\mathcal{P}^{O}=\mathcal{NP}^{O}$; thus, we have ${\rm co}\mathcal{P}^{O}={\rm co}\mathcal{NP}^{O}$. Clearly, in this scenario, to separate $\mathcal{NP}^{O}$ from ${\rm co}\mathcal{NP}^{O}$ is equivalent to separating $\mathcal{P}^{O}$ from ${\rm co}\mathcal{P}^{O}$. Moreover, if the set of all ${\rm co}\mathcal{P}^{O}$ machines is not enumerable, then of course we can not apply simulation techniques. Thus, we assume that the set of ${\rm co}\mathcal{P}^{O}$ machines is enumerable, as the premise of the theorem states. Since we have assumed that the set of all ${\rm co}\mathcal{P}^{O}$ machines is enumerable, the rest of the proof is similar to Remark 4.2. Namely, the simulation technique is unable to create a language not in ${\rm co}\mathcal{P}^{O}$ via a universal deterministic oracle Turing machine $U^{O}$, because a polynomial-time deterministic oracle Turing machine that works under the “there exists no accepting path” condition is still a polynomial-time deterministic oracle Turing machine. In short, the “there exists at least one accepting path” condition and the “there exists no accepting path” condition for deterministic oracle Turing machines are symmetrical. Specifically, for an arbitrary deterministic oracle Turing machine $M^{O}$, if $M^{O}$ accepts the input $w$, then there is only one accepting path, and if $M^{O}$ rejects the input $w$, then there is also only one rejecting path. This is different from nondeterministic oracle Turing machines. If we apply the same techniques, i.e., the simulation techniques, to create a language not in ${\rm co}\mathcal{P}^{O}$, then the universal deterministic oracle Turing machine $U^{O}$ (it runs within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$) needed to be constructed accepts the language $L(U^{O})\in\mathcal{P}^{O}$ (the proof is similar to that of Theorem 6.4 below). But, by the symmetrical properties between the “there exists at least one accepting path” condition and the “there exists no accepting path” condition with respect to deterministic oracle Turing machines, we clearly have that $L(U^{O})\in{\rm co}\mathcal{P}^{O}$. Thus, we are unable to claim that $L(U^{O})\notin{\rm co}\mathcal{P}^{O}$.
@@
    The simulation can be done within time
@@
Since the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines is enumerable, and apart from it, we further have the above rational assumptions, i.e., the following: (1) Any ${\rm co}\mathcal{NP}^{A}$ machine can be encoded to a string over $\{0,1\}$; (2) There exist universal nondeterministic oracle Turing machines with oracle $A$ that can simulate any other polynomial-time nondeterministic oracle Turing machine with oracle $A$, which is similar to the behavior of universal nondeterministic Turing machine $U$ constructed in Section 4; (3) The simulation can be done within time
@@
    $U^{A}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$ with oracle $A$ for some $j$ (by Definition 6.4, equivalently, $x$ can not be decoded into some polynomial-time nondeterministic oracle Turing machine $M^{A}_{j}$ for some $j$, i.e., $x$ represents the trivial oracle Turing machine with an empty next-move function), then GOTO $6$; else determine $t$, the number of tape symbols used by $M^{A}_{j}$; $s$, its number of states; and $k$, its order.  [^7] We suppose that the order $10^{k}1$ of $M^{A}_{j}$ for some $k\in\mathbb{N}_{1}$ is also encoded into the binary string representing ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{j}$, which is similar to Section 3. The third tape of $U^{A}$ can be used as “scratch” memory to calculate $t$.
@@
    By using its nondeterministic choices, $U^{A}$ moves as per the path given by $x$. The moves of $U^{A}$ are counted in binary in the block of tape $3$. If the counter on tape $3$ overflows, then $U^{A}$ halts. $U^{A}$ accepts $x$ if and only if there is at least one computation path from the start state of $U^{A}$ leading to the accept state, and the total number of moves can not exceed
@@
We claim that $L^{A}_{s}\notin{\rm co}\mathcal{NP}^{A}$. Indeed, suppose for the sake of contradiction that $L^{A}_{s}$ is decided by some $n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}^{A}$ machine $M^{A}_{i}$. Then, we may assume that $M^{A}_{i}$ is a two-tape (with an input tape [^8] Similarly, it can be proved that $t(n)$ time-bounded $k$-tape (additionally with an oracle tape) oracle nondeterministic Turing machines are equivalent to $t^{2}(n)$ time-bounded single-tape (additionally with an oracle tape) oracle nondeterministic Turing machines. and an oracle tape) $T(n)=n^{k}+k$ time-bounded ${\rm co}\mathcal{NP}^{A}$ machine. By Definition 6.4, we can regard it as an $n^{k}+k$ time-bounded nondeterministic oracle Turing machine with oracle $A$, and we can further let $U^{A}$ on input $1^{n-m}\langle M^{A}_{i}\rangle$ simulate $M^{A}_{i}$ on input $1^{n-m}\langle M^{A}_{i}\rangle$. Let $M^{A}_{i}$ have $s$ states and $t$ tape symbols, and the shortest length of $M^{A}_{i}$ is $m$, i.e.,
@@
If $\mathcal{P}^{A}\neq\mathcal{NP}^{A}$. Assume that the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines is enumerable. Further, under the above rational assumptions (i.e., Assumption 6.1 above), we have
@@
Under some rational assumptions (i.e., Assumption 6.1 above), and if $\mathcal{P}^{A}\neq\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, then the set $\mathcal{T}^{A}$ of all ${\rm co}\mathcal{NP}^{A}$ machines is not enumerable. Thereby, ordinary simulation techniques will generally not apply to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.
Now, we are naturally at the right point to present the proof of Theorem 1.3 as follows:
Proof of Theorem 1.3. Note that the statement of Theorem 1.3 is precisely the same as Theorem 6.6; therefore, we have done the proof.